OpenStudy (anonymous):

Find the LCM. t^3+8t^2+16, t^2-8t

OpenStudy (anonymous):

it sud be t^3 + 8t^2 + 16t

OpenStudy (turingtest):

assuming sheg is right (which makes sense) it should be t^3+8t^2+16t=t(t+4)^2 t^2-8t=t(t-4) LCM=t(t-4)(t+4)^2=(t^2-8t)(t^2+8t+16)=t^4+56t^2-128t